OpenStudy (ssprague_77):

If 9 is subtracted from two-thirds of a number, the result is 7. Find the number. @563blackghost

OpenStudy (caseysorensen):

(2/3 * x) - 9 = 7 2/3 * x = 7 + 9 = 16. Divide both sides by 2/3 x = 16 divided by 2/3 = 24.

OpenStudy (ssprague_77):

Thank you

OpenStudy (caseysorensen):

your welcome

alones (alones):

This question also belongs to math section.

OpenStudy (caseysorensen):

she already no that I already told her

OpenStudy (ssprague_77):

Yeah, I thought that I was in Mathematics, sorry I didn't notice I was in History instead.. I'm really sorry. I'll pay attention next time.

alones (alones):

Lol it happens cx

OpenStudy (caseysorensen):

lol k

OpenStudy (ssprague_77):

lol XD